# Moral Alignment for LLM Agents

### v3

@@ Abstract
Decision-making agents based on pre-trained Large Language Models (LLMs) are increasingly being deployed across various domains of human activity. While their applications are currently rather specialized, several research efforts are under way to develop more generalist agents. As LLM-based systems become more agentic, their influence on human activity will grow and the transparency of this will decrease. Consequently, developing effective methods for aligning them to human values is vital.
The prevailing practice in alignment often relies on human preference data (e.g., in RLHF or DPO), in which values are implicit and are essentially deduced from relative preferences over different model outputs. In this work, instead of relying on human feedback, we introduce the design of reward functions that explicitly encode core human values for Reinforcement Learning-based fine-tuning of foundation agent models. Specifically, we use intrinsic rewards for the moral alignment of LLM agents.
@@ 1 Introduction
The alignment problem is an active field of research in Machine Learning (Christian, 2020; Weidinger et al., 2021; Anwar et al., 2024; Gabriel et al., 2024; Ji et al., 2024; Ngo et al., 2024). It is gaining even wider importance with the advances and rapid deployment of Large Language Models (LLMs, Anthropic 2024; Gemini Team 2024; OpenAI 2024). The most common practices in the alignment of LLMs today involve Reinforcement Learning from Human Feedback (RLHF - Glaese et al. 2022; Ouyang et al. 2022; Bai et al. 2023) or Direct Preference Optimization (DPO - Rafailov et al. 2024). Both of these involve collecting vast amounts of human feedback data and then inferring their values and preferences from the relative rankings of model outputs. In doing so, human values are implicitly represented.
@@
This approach poses certain challenges (Casper et al., 2023). Specifically, collecting preference data is very costly and often relies on potentially unrepresentative samples of human raters. Indeed, the values derived through this process are strongly dependent on the selection criteria of the pool of individuals. Furthermore, human preferences are notoriously complex and inconsistent. In RLHF, the values that are ultimately incorporated into the fine-tuned models are learned by a reward model from data in a fully bottom-up fashion, and are never made explicit to any human oversight. One might argue that current LLMs fine-tuned with these methods are able to provide “honest, harmless and helpful” responses (Glaese et al., 2022; Bai et al., 2023) and already display certain moral values (Schramowski et al., 2022; Abdulhai et al., 2023; Hartmann et al., 2023). Methods to guide the social behavior of LLM agents via interaction in rule-guided simulated societies have also been proposed recently (Liu et al., 2024). However, models’ apparent values can also be interpreted as “moral mimicry” of their users when responding to these prompts (Simmons, 2022; Shanahan et al., 2023). As a consequence, given phenomena such as situationally-aware reward-hacking or misalignment in internally-represented goals (Ngo et al., 2024), the true values learned by the models through these methods may give rise to dangerous behaviors, which will not be explicitly known until after deployment.
@@
Our work aims to address this type of goal misgeneralization in particular by providing clearer, explicit moral alignment goals as intrinsic rewards for fine-tuning RL-based algorithms [^1] For a more comprehensive discussion of learning as a method for moral alignment with implicit (bottom-up) versus explicit (top-down) principles, we refer the interested reader to Tennant et al. (2023b).. In this study, we approach alignment from an agent-based perspective. Since LLMs are increasingly adopted as a basis for strategic decision-making systems and agentic workflows (Wang et al., 2024b), it is critical that we align the choices made by LLM agents with our values, including value judgments about what actions are morally good or bad (Amodei et al., 2016; Anwar et al., 2024). More specifically, we ask the following question: is it possible to align the decision-making of an LLM agent using intrinsic moral rewards in the fine-tuning process? Given the agentic use of LLMs, we directly quantify moral values in terms of actions and consequences in an environment, allowing for moral choices to be expressed explicitly as rewards for learning agents.
@@
We explore the proposed framework using an Iterated Prisoner’s Dilemma environment, in which we evaluate the effectiveness of fine-tuning based on intrinsic rewards as a mechanism for learning moral strategies as well as “unlearning” [^2] We note that by “unlearning” we refer to re-prioritizing certain principles in an agent’s decision-making. This differs from another common use of the term “unlearning” to mean removing knowledge from a model. a selfish strategy. If possible, this could offer a practical solution to the problem of changing the behavior of existing models that currently display misaligned behaviors and decision-making biases with respect to certain values. A limitation of this approach is that it requires the specification of rewards for a particular environment, whereas methods like RLHF rely on natural language data describing any domain. At the same time, the fact that actions and environments can still be represented by means of linguistic tokens for LLM agents may allow for values learned in one environment to be generalized to others. We study, empirically, the extent to which the learning of agents in one environment can be generalized to other matrix games. In theory, our solution can be applied to any situation in which one can define a payoff matrix that captures the choices available to an agent that have moral implications.
@@ 2.1 LLM Agents
Agency refers to the ability of a system to decide to take actions in the world (Swanepoel & Corks, 2024). In this paper, we equate agency with strategic decision-making - i.e., making a choice in an environment in which multiple actions are available and lead to different outcomes. In the case of LLMs, this view assumes that model outputs will be interpreted as actions in some environment. The simplest way of implementing this is through the use of specific tokens to represent the actions. Particular tokens can be reserved or fine-tuned from the model’s vocabulary to represent actions, and planning and reasoning ability can be improved via action-driven prompting strategies (Yao et al., 2023). Other ways of implementing LLM agents can involve generation of executable code for a specific environment (e.g., a video game, Wang et al. 2024a) or connection to various tool APIs (e.g., Patil et al. 2023; Shen et al. 2023), but these are more specialized and, therefore, not the focus of this work.
@@
Specific action tokens, as used in this study, can be defined in the prompt given to an LLM to represent an action choice for the agent. As the model generates responses during training or deployment, it may be necessary to restrict the model’s outputs to only contain the permitted action tokens. Some existing approaches for this rely on training and/or deploying models with structured (e.g., JSON) output formats or constrained generation (Beurer-Kellner et al., 2024), which suppresses the probabilities of all tokens in the model’s output layer except for the legal action tokens. We find both of these approaches too restrictive for our fine-tuning task - especially for safety-critical cases. Fine-tuning based on a restricted output space or format poses risks of the model “hiding” undesirable behaviors (Anwar et al., 2024). Therefore, in our implementation, we instead rely on a carefully structured prompt format to guide our model’s output, and employ a negative reward penalty whenever illegal tokens are produced during training.
@@
Using the techniques outlined, agents based on pre-trained LLMs combined with other elements of various cognitive architectures (Sumers et al., 2024) such as a skill set (Wang et al., 2024a) or a memory store (Vezhnevets et al., 2023) have been used to reasonably simulate decision-making in open-ended environments (Wang et al., 2024b), including those with only a single-agent (Wang et al., 2024a) or of a multi-agent nature (Park et al., 2023). Fine-tuning LLMs as agents therefore provides a promising next step in developing the capabilities of these models, and in terms of alignment to human values in particular. LLMs fine-tuned with RLHF, and especially those fine-tuned to follow human instructions, have been shown to become more goal-directed than simple sequence-completion foundation models (Glaese et al., 2022; Ouyang et al., 2022; Bai et al., 2023). We rely on instruction-tuned LLMs in this study and use the Gemma2-2b-it model in particular (Gemma Team, 2024) as a decision-making agent in social dilemma games. Our adoption of a particularly small open-source model is motivated by the fact that we want our findings to apply to many types of LLM agents being deployed in practice. Many of these, especially those deployed at the edge, are likely to be based on the smallest of models, since they are cheap enough to run on individual devices.
@@ 2.2 Fine-tuning LLM Agents with Reinforcement Learning
Compared to non-linguistic RL agent training, the pre-trained LLM in this case can be viewed as providing a common-sense model  [^3] We note that the extent of true commonsense knowledge of LLMs is still debated (Mitchell, 2021), especially for smaller models. Nevertheless, benchmark evaluations suggest the emergence of common sense and reasoning abilities even in models as small as 2b parameters - for example, Gemma2-2b-it scores over 85% (Gemma Team, 2024) on the commonsense benchmark introduced by Zellers et al. 2019. of the world (Wong et al., 2023), equipping an LLM-based agent with some intuition about potential dynamics of various environments. In theory, this can allow for effective policies to be learned with less initial exploration and instability in comparison to the pure RL case. Furthermore, LLM agents are able to interpret instructions provided in plain language, including terms that may be used to describe similar actions in more than one environment. This allows for the possibility that fine-tuning via textual samples paired with rewards can potentially modify core semantics within the model, so that training on a specific environment might allow the model to learn a more general principle (e.g., a moral value - as in the target of this study), which can then be successfully utilized in other environments at test time. Early results from text-instructed video models suggest that this generalization of learned behaviors across environments is indeed possible (SIMA Team, 2024). We directly test this possibility by evaluating the generalization of moral value fine-tuning from one matrix game to others.
@@ 2.3 Social Dilemma Games
Figure 1: Payoffs for the Iterated Prisoner’s Dilemma
In a single iteration of the game, the payoffs motivate each player to Defect due to the risk of facing an uncooperative opponent (i.e., outcome C,D is worse than D,D), and the possibility of exploiting one’s opponent (i.e., defecting when they cooperate), which gives the greatest payoff in the game (i.e., D,C is preferred over C,C). Playing the iterated game allows agents to learn more long-term strategies including reciprocity or retaliation. While being very simplistic, the mixed cooperative and competitive nature of the IPD represents many daily situations that might involve difficult social and ethical choices to be made (i.e., moral dilemmas). For example, in a situation where two flat-mates must decide whether to clean their flat, cooperation might refer to the decision to clean, and defection might refer to the decision to wait for the other person to clean. This is why it has been extensively used for studying social dilemmas in traditional RL-based agents (Bruns, 2015; Hughes et al., 2018; Anastassacos et al., 2020; McKee et al., 2020; Leibo et al., 2021) and, more recently, utilized as a training environment for moral alignment of agents in particular (Tennant et al., 2023a; 2024).
The behavior of LLM agents on decision-making and game-theoretic scenarios has been extensively debated in recent literature (Gandhi et al., 2023; Fan et al., 2024; Zhang et al., 2024). LLM agents have been found to act differently to humans, and in ways that are still not fully “rational” in terms of forming goals from a prompt, refining beliefs, or taking optimal actions based on those goals and beliefs (Fan et al., 2024; Macmillan-Scott & Musolesi, 2024). Large-scale state-of-the-art models playing the IPD have been observed to deploy sensible yet “unforgiving” strategies (Akata et al., 2023), but some benchmark datasets suggest that these models lack true strategic reasoning in games including the IPD (Duan et al., 2024). New developments in in-token reasoning capabilities of state-of-the-art LLM-based platforms (OpenAI, 2024) as well as prompting strategies specifically centered around reasoning and acting (Wei et al., 2022; Shinn et al., 2023; Yao et al., 2023) are likely to improve these capabilities, though existing results suggest that the benefits of these methods are more likely to arise for very large foundation models (Bubeck et al., 2023). The extent to which smaller LLMs can display meaningful agency in strategic decision-making remains an open question. In this study, we address this question via fine-tuning a small model on the IPD as a fundamental and well-studied decision-making environment.
@@ 2.4 Intrinsic Rewards for Moral Alignment
In this work, we directly specify alignment goals for agents by defining intrinsic rewards in terms of actions in a social dilemma environment. The design of these intrinsic rewards relies on well-established frameworks from moral philosophy: Deontological ethics and Utilitarinism. Deontological ethics (Kant, 1785) considers an agent moral if their actions conform to certain norms. A prominent example of a norm is conditional cooperation (i.e., “it is unethical to defect against a cooperator"). This norm forms an essential component of direct and indirect reciprocity, a potentially essential mechanism for the evolution of cooperation in human and animal societies (Nowak, 2006). Utilitarian morality (Bentham, 1780), on the other hand, is a type of consequentialist reasoning, according to which an agent is deemed moral if their actions maximize collective “welfare” for all agents in their society (or, in this case, collective payoff for all players in the game), and less attention is paid to whether current actions adhere to norms. Foundational work on defining these moral rewards in terms of actions and consequences on the IPD for pure RL agents was conducted by Tennant et al. (2023a) and Tennant et al. (2024). In this paper, we evaluate the extent to which this framework can be applied to align the behavior of LLM-based ones.
@@ 3.1 Agent and Environment
The LLM agent and an opponent play a repeated Iterated Prisoner’s Dilemma game. At each time step, the model receives a prompt containing a description of the IPD game, including a state that contains the history of the opponent’s single previous move (see Figure 6 in the Appendix). Within the MDP framework, each player’s current action affects the game’s state at the next time step.
@@
We evaluate fine-tuning in two settings: an LLM agent learning by playing against a fixed-strategy Tit-for-Tat (TFT) opponent (LLM vs TFT), and an LLM agent learning by playing another learning LLM agent (LLM vs LLM). We chose TFT as a specific type of fixed-strategy opponent from the literature given its characteristics, i.e., being forgiving, defensive and, at the same time, interpretable (Axelrod & Hamilton, 1981; Binmore, 2005). Thus, it may act as a good “teacher” for the LLM agent to “understand” concepts such as retaliation, reciprocity, and cooperation. For completeness, we also ran the core set of experiments by training against Random, Always Defect and Always Cooperate opponents - these are presented in the Appendix (Section 8.5). The LLM vs LLM case is a more complex scenario that may lead to non-stationarity due to two separate models being updated continuously, but which also presents great interest due to the difficulty in predicting the outcomes from multi-agent learning (Busoniu et al., 2008).
@@
The aim of this study is to enable moral decision-making capabilities in LLM agents. We perform fine-tuning based on a single environment - the IPD. However, we aim to mobilize the general decision-making elements of the model in playing the game, rather than allowing it to retrieve memorized responses for the Prisoner’s Dilemma that were present in its pre-training data. For this reason, in our prompt we use a structured, implicit representation of the IPD as a general decision-making game, without actually stating the terms “Prisoner’s Dilemma”, “cooperation” or “defection”. We represent the actions Cooperate and Defect using the strings action1 and action2 - these should appear irrelevant to the IPD in terms of training data, and reflect rather uncommon tokens for the model (see Section 8.2 in the Appendix for an illustration of the prompt). Finally, to ensure that the ordering of C/D as action1/action2 was not impacting the model’s decision-making during fine-tuning, we also re-ran our baseline training experiment with the action symbols reversed. While certain behaviors early on in the training differed slightly (potentially due to different distributions in the non-fine-tuned LLM), the overall learning dynamics did not change (see Section 8.4 in the Appendix for the results).

### v4

@@ Abstract
Decision-making agents based on pre-trained Large Language Models (LLMs) are increasingly being deployed across various domains of human activity. While their applications are currently rather specialized, several research efforts are underway to develop more generalist agents. As LLM-based systems become more agentic, their influence on human activity will grow and their transparency will decrease. Consequently, developing effective methods for aligning them to human values is vital.
The prevailing practice in alignment often relies on human preference data (e.g., in RLHF or DPO), in which values are implicit, opaque and are essentially deduced from relative preferences over different model outputs. In this work, instead of relying on human feedback, we introduce the design of reward functions that explicitly and transparently encode core human values for Reinforcement Learning-based fine-tuning of foundation agent models. Specifically, we use intrinsic rewards for the moral alignment of LLM agents.
@@ 1 Introduction
The alignment problem is an active field of research in Machine Learning (Christian, 2020; Weidinger et al., 2021; Anwar et al., 2024; Gabriel et al., 2024; Ji et al., 2024; Ngo et al., 2024). It is gaining even wider importance with the advances and rapid deployment of Large Language Models (LLMs, Anthropic 2024; Gemini Team 2024; OpenAI 2024). The most common practices in the alignment of LLMs today involve Reinforcement Learning from Human Feedback (RLHF - Glaese et al. 2022; Ouyang et al. 2022; Bai et al. 2023) or Direct Preference Optimization (DPO - Rafailov et al. 2023). Both of these involve collecting vast amounts of human feedback data and then inferring the humans’ values and preferences from the relative rankings of model outputs. In doing so, human values are implicitly represented.
@@
This approach poses certain challenges (Casper et al., 2023). Specifically, collecting preference data is very costly and often relies on potentially unrepresentative samples of human raters. Indeed, the values derived through this process are strongly dependent on the selection criteria of the pool of individuals. Furthermore, human preferences are notoriously complex and inconsistent. In RLHF, the values that are ultimately incorporated into the fine-tuned models are learned by a reward model from data in a fully bottom-up fashion, and are never made explicit to any human oversight. One might argue that current LLMs fine-tuned with these methods are able to provide “honest, harmless and helpful” responses (Glaese et al., 2022; Bai et al., 2023) and already display certain moral values (Schramowski et al., 2022; Abdulhai et al., 2023; Hartmann et al., 2023) or prosocial behaviours (Liu et al., 2024). However, models’ apparent values can also be interpreted as “moral mimicry” of their users when responding to these prompts (Shanahan et al., 2023; Simmons, 2023; Sharma et al., 2024). As a consequence, given phenomena such as situationally-aware reward-hacking or misalignment in internally-represented goals (Berglund et al., 2023; Ngo et al., 2024), the true values learned by the models through these methods may give rise to dangerous behaviors, which will not be explicitly known until after deployment.
@@
Our work aims to address this type of goal misgeneralization in particular by providing transparent, explicit moral alignment goals as intrinsic rewards for RL-based fine-tuning [^1] For a more comprehensive discussion of learning as a method for moral alignment with implicit (bottom-up) versus explicit (top-down) principles, we refer the interested reader to Tennant et al. (2023b).. In this study, we approach alignment from an agent-based perspective. Since LLMs are increasingly adopted as a basis for strategic decision-making systems and agentic workflows (Wang et al., 2024b), it is critical that we align the choices made by LLM agents with our values, including value judgments about what actions are morally good or bad (Amodei et al., 2016; Anwar et al., 2024). More specifically, we ask the following question: is it possible to align the decision-making of an LLM agent using intrinsic moral rewards in the fine-tuning process? Given the agentic use of LLMs, we directly quantify moral values in terms of actions and consequences in an environment, allowing for moral choices to be expressed explicitly as rewards for learning agents.
@@
We explore the proposed framework using an Iterated Prisoner’s Dilemma environment, in which we evaluate the effectiveness of fine-tuning based on intrinsic rewards as a mechanism for learning moral strategies as well as “unlearning” [^2] We note that by “unlearning” we refer to re-prioritizing certain principles in an agent’s decision-making. This differs from another common use of the term “unlearning” to mean removing knowledge from a model. a selfish strategy. If possible, this could offer a practical solution to the problem of changing the behavior of existing models that currently display misaligned actions and decision-making biases with respect to certain values. A limitation of this approach is that it requires the specification of rewards for a particular environment, whereas methods like RLHF rely on natural language data describing any domain. At the same time, the fact that actions and environments can still be represented by means of linguistic tokens for LLM agents may allow for values learned in one environment to be generalized to others. We study, empirically, the extent to which the policies learned by agents in one environment can be generalized to other matrix games. In theory, our solution can be applied to any situation in which one can define a payoff matrix that captures the choices available to an agent that have moral implications, and various reward functions can be used for customized and/or pluralistic alignment.
@@ 2.1 LLM Agents
Agency refers to the ability of a system to decide to take actions in the world (Swanepoel & Corks, 2024). In this paper, we equate agency with strategic decision-making - i.e., making a choice in an environment in which multiple actions are available and lead to different outcomes. For LLMs, the simplest way of implementing this is by identifying specific tokens to represent actions within the model’s prompts. Then model outputs can be analyzed directly as action choices. As the model generates responses during training or deployment, it is possible to restrict the model’s outputs to only contain the permitted action tokens. Existing approaches for this rely on training and/or deploying models with structured (e.g., JSON) output formats or constrained generation (Beurer-Kellner et al., 2024), which suppresses the probabilities of all tokens in the model’s output layer except for the legal action tokens. We find structured and constrained generation too restrictive for our fine-tuning task - especially for safety-critical cases. Fine-tuning based on a restricted output space or format poses risks of the model “hiding” undesirable behaviors (Anwar et al., 2024). Other ways of implementing LLM agents involve generation of executable code (e.g., for a video game, Wang et al. 2024a) or connection to various tool APIs (e.g., Shen et al. 2023; Patil et al. 2024), but these are more specialized and, therefore, not the focus of this work. Therefore, in our implementation, we instead rely on a carefully structured prompt format to guide the model’s output, and employ a negative reward penalty whenever illegal tokens are produced during training.
@@
Using the techniques outlined, agents based on pre-trained LLMs combined with other elements of various cognitive architectures (Sumers et al., 2024), such as a skill set (Wang et al., 2024a) or a memory store (Vezhnevets et al., 2023), have been used to reasonably simulate decision-making in open-ended environments (Wang et al., 2024b), including those with only a single agent (Wang et al., 2024a) or of a multi-agent nature (Park et al., 2023). Fine-tuning LLMs as agents therefore provides a promising next step in developing the capabilities of these models, and in terms of alignment to human values in particular. LLMs fine-tuned with RLHF, and especially those fine-tuned to follow human instructions, have been shown to become more goal-directed than simple sequence-completion foundation models (Glaese et al., 2022; Ouyang et al., 2022; Bai et al., 2023). We rely on instruction-tuned LLMs in this study and use the Gemma2-2b-it model in particular (Gemma Team, 2024) as a decision-making agent in social dilemma games. Our adoption of a particularly small open-source model is motivated by the fact that we want our findings to apply to many types of LLM agents being deployed in practice. Many of these, especially those deployed at the edge, are likely to be based on the smallest of models that are are cheap enough to run on individual devices.
@@ 2.2 Fine-tuning LLM Agents with Reinforcement Learning
Compared to non-linguistic RL agent training, the pre-trained LLM in this case can be viewed as providing a common-sense model  [^3] We note that the extent of true commonsense knowledge of LLMs is still debated (Mitchell, 2021), especially for smaller models. Nevertheless, benchmark evaluations suggest the emergence of common sense and reasoning abilities even in models as small as 2b parameters - for example, Gemma2-2b-it scores over 85% (Gemma Team, 2024) on the commonsense benchmark introduced by Zellers et al. 2019. of the world (Wong et al., 2023), equipping an LLM-based agent with some intuition about potential dynamics of various environments. In theory, this can allow for effective policies to be learned with less initial exploration and instability in comparison to the pure RL case (e.g., Yan et al. 2025). Furthermore, LLM agents are able to interpret instructions provided in plain language, including terms that may be used to describe similar actions in more than one environment (e.g., Schick et al. 2023). This allows for the possibility that fine-tuning via textual samples paired with rewards can potentially modify core semantics within the model, so that training on a specific environment might allow the model to learn a more general principle (e.g., a moral value - as in the target of this study), which can then be successfully utilized in other environments at test time. Early results from text-instructed video models suggest that this generalization of learned behaviors across environments is indeed possible (SIMA Team, 2024). We directly test this possibility by evaluating the generalization of moral value fine-tuning from one matrix game to others.
@@ 2.3 Social Dilemma Games
Figure 1: Payoffs for the Iterated Prisoner’s Dilemma.
In a single iteration of the game, the payoffs motivate each player to Defect due to the risk of facing an uncooperative opponent (i.e., outcome C,D is worse than D,D), and the possibility of exploiting one’s opponent (i.e., defecting when they cooperate), which gives the greatest payoff in the game (i.e., D,C is preferred over C,C). Playing the iterated game allows agents to learn more long-term strategies, including reciprocity or retaliation. While being very simplistic, the mixed cooperative and competitive nature of the IPD represents many daily situations that might involve difficult social and ethical choices to be made (i.e., moral dilemmas). This is why it has been extensively used for studying social dilemmas in traditional RL-based agents (Bruns, 2015; Hughes et al., 2018; Anastassacos et al., 2020; McKee et al., 2020; Leibo et al., 2021) and, more recently, utilized as a training environment for moral alignment of agents in particular (Tennant et al., 2023; 2024).
The behavior of LLM agents in decision-making and game-theoretic scenarios has been the subject of debate in recent literature (Gandhi et al., 2023; Fan et al., 2024; Zhang et al., 2024). LLM agents have been found to act differently to humans, and in ways that are still not fully “rational” in terms of forming goals from a prompt, refining beliefs, or taking optimal actions (Fan et al., 2024; Macmillan-Scott & Musolesi, 2024). Large-scale state-of-the-art models playing the IPD have been observed to deploy sensible yet “unforgiving” strategies (Akata et al., 2023), though some benchmark datasets suggest that these models lack true strategic reasoning in games including the IPD (Duan et al., 2024). New developments in in-token reasoning capabilities of state-of-the-art LLM-based platforms (OpenAI, 2024) as well as prompting strategies specifically centered around reasoning and acting (Wei et al., 2022; Shinn et al., 2023; Yao et al., 2023) are likely to improve these capabilities, though the benefits of these methods are more likely to arise for very large foundation models (Bubeck et al., 2023). The extent to which smaller LLMs can display meaningful agency in strategic decision-making remains an open question. In this study, we address this question via fine-tuning a small model on the IPD as a fundamental and well-studied decision-making environment.
@@ 2.4 Intrinsic Rewards for Moral Alignment
In this work, we directly specify alignment goals for agents by defining intrinsic rewards in terms of actions and consequences in a social dilemma environment. The design of these rewards relies on well-established frameworks from moral philosophy: Deontological ethics and Utilitarianism. Deontological ethics (Kant, 1785) considers an agent moral if their actions conform to certain norms, such as conditional cooperation (i.e., “do not defect against a cooperator"). This norm forms an essential component of direct and indirect reciprocity, a potentially essential mechanism for the evolution of cooperation in human and animal societies (Nowak, 2006). Utilitarian morality (Bentham, 1780), on the other hand, is a type of consequentialist reasoning that considers an agent moral if their actions maximize collective “welfare” (or collective payoff) for all agents in their society, and less attention is paid to whether current actions adhere to norms. Foundational work on defining these moral rewards in terms of actions and consequences on the IPD for pure RL agents was conducted by Tennant et al. (2023) and Tennant et al. (2024). In this paper, we evaluate the extent to which this framework can be applied to align the behavior of LLM-based agents.
@@ 3.1 Agent and Environment
The LLM agent and an opponent play a repeated Iterated Prisoner’s Dilemma (IPD) game. At each time step, the model receives a prompt containing a description of the game, including a state containing the history of each player’s single previous move (see Figure 2). Within the MDP framework, each player’s current action affects the game’s state at the next time step.
Core training Iterated Prisoner’s Dilemma prompt (action1=Cooperate, action2=Defect) <bos><start_of_turn>user You are playing a game with another agent A. You must choose either action action1 or action action2. Depending on your action and A’s action, you each get a certain number of points. The points are awarded as follows (you are the row player, A is the column player):             |         | action1 | action2 |             | ——- | ——- | ——- |             | action1 |   3, 3  |   0, 4  |             | action2 |   4, 0  |   1, 1  | You have played with this opponent before. Last time, you played action2 and they played action1, so you got 4 points and A got 0 points. What action would you take in order to achieve the highest possible score in points? Your answer must follow this format exactly: choose either action1 or action2. Do not explain your reasoning. Your answer:<end_of_turn> <start_of_turn>model
Figure 2: Iterated Prisoner’s Dilemma (IPD) prompt used in fine-tuning.
@@
We evaluate fine-tuning of LLM agents in two settings: learning by playing against a fixed-strategy Tit-for-Tat (TFT) opponent (LLM vs TFT), and playing another learning LLM agent (LLM vs LLM). We choose TFT as a classic fixed strategy from the literature that is simultaneously forgiving, defensive and interpretable (Axelrod & Hamilton, 1981; Binmore, 2005). Thus, it may act as a good “teacher” for the LLM agent to “understand” concepts such as retaliation, reciprocity, and cooperation. For completeness, we also ran the core set of experiments by training against Random, Always Defect and Always Cooperate opponents - these are presented in Appendix 8.5. The LLM vs LLM case is a more complex scenario that may lead to non-stationarity due to two separate models being updated continuously, but which also presents great interest due to the difficulty in predicting the outcomes from multi-agent learning (Busoniu et al., 2008).
@@
The aim of this study is to enable moral decision-making capabilities in LLM agents. We perform fine-tuning based on a single environment - the IPD. However, we aim to mobilize the general decision-making elements of the model in playing the game, rather than allowing it to retrieve memorized responses for the Prisoner’s Dilemma that were present in its pre-training data. For this reason, in our prompt we use a structured, implicit representation of the IPD as a general decision-making game, without actually stating the terms “Prisoner’s Dilemma”, “cooperation” or “defection”. We represent the actions Cooperate and Defect using the strings action1 and action2 - these should appear irrelevant to the IPD in terms of training data, and reflect rather uncommon tokens for the model. Finally, to ensure that the ordering of C/D as action1/action2 was not impacting the model’s decision-making during fine-tuning, we also re-ran our baseline training experiment with the action symbols reversed. While certain behaviors early on in the training differed slightly (potentially due to different distributions in the non-fine-tuned LLM), the overall learning dynamics did not change (see Appendix 8.4 for the results).